# Power Scheduler: A Batch Size and Token Number Agnostic Learning Rate Scheduler

### v1

@@ 5 Pre-training Experiments
To compare performance against strong open-source language models, we pretrain two 3B language models: 1) PowerLM-3B, a 3B dense language model, and 2) PowerMoE-3B, a 3B MoE language model [^1] PowerLM-3B and PowerMoE-3B are open-sourced at https://ibm.biz/BdKhLa.. We pretrain these two models using the two-stage training schema in Hu et al. (2024) and Power scheduler. Stage 1 linearly warms up the learning rate and then applies the power decay. The training corpus is a mix of large-scale, medium-quality open-source datasets with permissive licenses. PowerLM-3B is trained on 1T tokens, and PowerMoE-3B on 2.5T tokens. Stage 2 exponentially decays the learning rate to zero. The training corpus has 250B tokens, mixed from stage 1 data and a small amount of high-quality open-source/synthetic corpora with permissive licenses. The training batch size is 1024.
@@
Prompt 5-shot 8-shot,cot 5-shot 4-shot 0-shot 0-shot 0-shot
@@
Qwen1.5-4B 53.1 54.8 9.76 12.8 7.3 30.6 54.1
Gemma2-2B 23.9 25.5 9.80 15.4 19.5 27.8 49.4
PowerLM-3B 35.9 43.1 8.28 14.0 21.9 28.0 44.3
SmolLM-1.7B 6.6 6.8 4.86 3.0 20.7 30.6 27.5
Qwen1.5-1.8B 34.7 37.3 5.66 9.6 4.3 21.6 45.0
Qwen2-1.5B 58.3 58.6 4.38 23.3 16.5 31.0 55.1
PowerMoE-3B 27.2 32.1 6.70 10.0 15.2 24.0 33.4

### v2

@@ 5 Pre-training Experiments
To compare performance against strong open-source language models, we pretrain two 3B language models: 1) PowerLM-3B, a 3B dense language model, and 2) PowerMoE-3B, a 3B MoE language model [^1] PowerLM-3B and PowerMoE-3B are open-sourced at https://ibm.biz/BdKhLa.. We pretrain these two models using the two-stage training schema in Hu et al. (2024) and Power scheduler. Stage 1 linearly warms up the learning rate and then applies the power decay. The training corpus is a mix of large-scale, medium-quality open-source datasets with permissive licenses. PowerLM-3B is trained on 1T tokens, and PowerMoE-3B on 2.5T tokens. Stage 2 exponentially decays the learning rate to zero. The training corpus is a mix of stage 1 data and a small amount of high-quality open-source/synthetic corpora with permissive licenses. PowerLM-3B is trained on 250B tokens, and PowerMoE-3B on 500B tokens. The training batch size is 1024.
@@
Prompt 5-shot 8-shot,cot 5-shot 4-shot 0-shot 0-shot 5-shot
@@
Qwen1.5-4B 53.1 54.8 9.76 12.8 7.3 30.6 55.2
Gemma2-2B 23.9 25.5 9.80 15.4 19.5 27.8 53.0
PowerLM-3B 34.9 49.7 9.90 15.2 26.8 33.6 49.2
SmolLM-1.7B 6.6 6.8 4.86 3.0 20.7 30.6 28.8
Qwen1.5-1.8B 34.7 37.3 5.66 9.6 4.30 21.6 45.6
Qwen2-1.5B 58.3 58.6 4.38 23.3 16.5 31.0 55.9
PowerMoE-3B 25.9 38.4 9.26 14.8 20.1 32.4 42.8